# The reflection complexity of sequences over finite alphabets

### v1

@@ 4.5 Complementation-symmetric Rote sequences
In fact, with each Sturmian sequence $\mathbf{s}$, there are two associated complem-entation-symmetric Rote sequences $\mathbf{x}$ and $\mathbf{x}^{\prime}$ with $\mathbf{x}^{\prime}=E(\mathbf{x})$. The factors in $\mathbf{s}$ and its corresponding Rote sequences are closely related as shown below.
Proposition 49 ([78]; also see [62, Proposition 2] or [63, Lemma 2.7]).
Let $\mathbf{s}$ be a Sturmian sequence and let $\mathbf{x}$ be the complementation-symmetric Rote sequence such that $\mathbf{s}=\Delta(\mathbf{x})$. Then $u$ is a factor of $\mathbf{s}$ if and only if both words $v,v^{\prime}$ such that $u=\Delta(v)=\Delta(v^{\prime})$ are factors of $\mathbf{x}$. Furthermore, for every $n\geq 0$, $x$ occurs at position $n$ in $\mathbf{s}$ if and only if $v$ or $v^{\prime}$ occurs at position $n$ in $\mathbf{x}$.
@@
Let $\mathbf{x}$ be a complementation-symmetric Rote sequence. We clearly have $r_{\mathbf{x}}(0)=1$. Now, for $n\geq 1$[4, Theorem 8] states that $\operatorname{Pal}_{\mathbf{x}}(n)=2$. We finish the proof using Lemma 50 and Theorems 9(b). ∎
5 Reversal-closed and rich sequences
Proposition 52.
Proof.
It is enough to combine Remark 10 and Theorems 17(b) and 9(b). ∎
Rich sequences have several equivalent definitions. It is known that a word $w$ contain at most $|w|+1$ many palindromic factors [42]. A sequence is called rich if each factor contains the maximal number of palindromic factors.
@@ 5 Reversal-closed and rich sequences
From [24, Theorem 1.1], the sequence $\mathbf{x}$ is rich if and only if the inequality in Theorem 17(b) is an equality. The result then follows from Theorem 9(b). ∎
Those among binary quasi-Sturmian sequences that are coding of rotations are rich, see [20, Thm. 19].
@@
Using Theorem 9(b) and [80, Cor. 2.27 and 2.29], it is possible to bound the reflection complexity of rich sequences as follows.
@@
Other sequences have a reflection complexity satisfying the equality of Theorem 53. For instance, it is the case of complementation–symmetric sequences, sequences canonically associated with some specific Parry numbers, and sequences coding particular interval exchange transformations. For more details, see [14, Section 3].
@@ 6 Automatic sequences
In this section, we study the reflection complexity of automatic sequences. First, in a positional numeration system $U$ having an adder, we show that if a sequence is $U$-automatic, then its reflection complexity is a $U$-regular sequence. Furthermore we show how to effectively compute a linear representation for the sequence, making use of the free software Walnut [68, 83]. Next, we explore the reflection complexity of some famous automatic sequences, namely the Thue–Morse, the period-doubling, generalized paperfolding, generalized Golay–Shapiro, and the Baum-Sweet sequences.
@@ 6.1 Reflection complexity is computably regular
We now show that the reflection complexity of any automatic sequence is regular.
@@
Let $U=(U(n))_{n\geq 0}$ be a positional numeration system such that there is an adder, i.e., addition is recognizable by an automaton reading $U$-representations, and let $\mathbf{x}$ be a $U$-automatic sequence. Then $(r_{\mathbf{x}}(n))_{n\geq 0}$ is a $U$-regular sequence. Furthermore, a linear representation for $(r_{\mathbf{x}}(n))_{n\geq 0}$ is computable from the DFAO for $\mathbf{x}$.
@@
6.2 The Thue–Morse and period-doubling sequences
We can compute a linear representation for the reflection complexity $r_{\mathbf{t}}(n)$ of the $2$-automatic Thue–Morse sequence, using the same approach as in the preceding section. Here we use the following Walnut code:
@@ 6.2 The Thue–Morse and period-doubling sequences
    There is an automaton of 14 states computing the first difference -rt(+n1)rt(n).
@@
    We can compute a linear representation for the first difference -rt(+n1)rt(n), and then use the “semigroup trick” [83, §4.11] to prove that the difference is bounded and find the automaton for it. It is displayed in Figure 1.
    Figure 1: Automaton computing -rt(+n1)rt(n) where t is the Thue–Morse sequence.
@@
The period-doubling sequence (d(n))n0 is a natural companion of the Thue–Morse sequence. Recall that t is the fixed point, starting with 0, of the morphism defined by 001 and 110. We similarly define p as the fixed point of the morphism 001 and 100. This gives us that p is 2-automatic as well. By defining d(n) as the highest power of 2, modulo 2, dividing +n1, the sequence p can be equivalently defined as the sequence (d(n))n0. The close relationship between t and p is captured by the identity =ρp(n)ρt(+n1)2 for all n. We may devise a close analogue of Theorem 57 for the reflection complexity of p, again with the use of Walnut. Explicitly, it can be shown that: For all n0, we have =rp(+2n1)+ρp(n)1, and, for all n2, we have
@@
and we may similarly devise an analogue of part (c) of Theorem 57. Observe that
=lim infnrp(n)n34and=lim supnrp(n)n56,
and similarly for the reflection complexity of t.
6.3 The generalized paperfolding sequences
@@
Now, by a known result on the appearance function of paperfolding sequences [83, Theorem 12.2.1], we know that every length-14 factor of a paperfolding sequence pf appears in a prefix of length 109, which is in turn specified by the first 7 unfolding instructions. We can then simply examine each of the 56 length-14 factors of these 128 (finite) words and verify that no factor is reflected. ∎
We can now prove the following result.
Theorem 59.
Let pf be a paperfolding sequence. Then
  • (a)
    For all n13, we have rf(n)=ρf(n)=4n.
  • (b)
    The reflection complexity of every paperfolding sequence is the same, and takes the values 2,3,6,7,12,15,22,24,32,36,42,46 for 1n12.
Proof.
We prove each item separately.
  • (a)
    For n14, the result follows from combining the results of Allouche [2] and Proposition 58. For =n13, we can verify the claim by explicit enumeration.
  • (b)
    The result for n13 follows from (a). For <n13 the result can be verified by enumeration of all length-109 prefixes of paperfolding sequences specified by instructions of length 7.
This ends the proof. ∎
6.4 The generalized Golay–Shapiro sequences
A generalized Golay–Shapiro sequence g is defined by taking the running sum, modulo 2, of a paperfolding sequence pf. The famous Golay–Shapiro sequence (also called the Rudin–Shapiro sequence) [51, 52, 79, 84] corresponds to the case of unfolding instructions 0(01)ω [4, Def. 6]. Note that the 2-automaticity of a generalized Golay–Shapiro sequence follows from that of its corresponding generalized paperfolding sequence.
We can prove the analogue of Proposition 58.
@@ 6.4 The generalized Golay–Shapiro sequences
Proof.
As above, it suffices to show that no Golay–Shapiro sequence contains a reflected factor of length 15.
Now, by a known result on the recurrence function of generalized Golay–Shapiro sequences [5, Proposition 4.1], we know that every length-15 factor of a paperfolding sequence pf appears in a prefix of length 2408, which is in turn specified by the first 12 unfolding instructions. We can then simply examine each of the 60 length-15 factors of these 4096 (finite) words and verify that no factor is reflected. ∎
@@
    The reflection complexity of every generalized Golay–Shapiro sequence is the same, and takes the values 2, 3, 6, 10, 14, 22, 30, 42, 48, 62, 72, 83, 92, 103 for 1n14.
@@
    The result for n15 follows from (a). For <n15 the result can be verified by enumeration of all length-2408 prefixes of paperfolding sequences specified by instructions of length 12.
@@ 6.5 The Baum–Sweet sequence
=b1101100101001001100100000100100101001001
be defined by =b(0)1 and for n1, b(n) is 1 if the base-2 expansion of n contains no block of successive zeros of odd length and 0 otherwise. It is 2-automatic as well. The factor complexity function for b is such that
@@
and the reflection complexity function for b starts with
@@
We can again compute a linear representation for rb(n) using the following Walnut code:
@@
Let b be the Baum-Sweet sequence. Then the first difference of the sequence rb(n) is 2-automatic, over the alphabet {1,2,,8}.
@@ 7 Further directions
This sequence is not automatic in a given so-called addable numeration system (where there is an adder). Therefore, we cannot use Walnut, in this case. However, an inductive argument can be applied to prove that =rc(n)ρc(n) for all n13.
Question 63.
@@
What is the reflection complexity of the Thue–Morse sequence over polynomial extractions, with regard to the work of Yossi Moshe [67]?
Question 66.
How could the upper bound in Theorem 20 be improved? If rx(n) is of the form Ω(n), then how could this be improved?
Question 67.
How does the reflection complexity compare with other complexity functions, as in the complexity functions listed in Section 1?
This leads us to ask about the respective growths of the complexities listed in Section 1, in particular for morphic sequences. In this direction, recall that the factor complexity of a morphic sequence is either Θ(1), Θ(n), Θ(nloglogn), Θ(nlogn) or Θ(n2), see [70] (more details can be found, e.g., in [35]; also see [41]). As an illustration with a result that has not been already cited above, a comparison between growths for the factor complexity and the Lempel-Ziv complexity can be found in [36]. We end with an easy result for the growth of the reflection complexity in the case of morphic sequences.
@@
The reflection complexity of a morphic sequence is either Θ(1), Θ(n), Θ(nloglogn), Θ(nlogn) or Θ(n2).
@@
Use the inequalities in Theorem 9: for any sequence x and for all n0, we have 12ρx(n)rx(n)ρx(n). ∎
@@ Acknowledgments
We thank Boris Adamczewski for discussions. John Campbell is grateful to acknowledge support from a Killam Postdoctoral Fellowship from the Killam Trusts. Manon Stipulanti is an FNRS Research Associate supported by the Research grant 1.C.104.24F. The research of Jeffrey Shallit is supported by NSERC grant 2018-04118.

### v4

@@ 6 Sturmian sequences and generalizations
In fact, for each Sturmian sequence $\mathbf{s}$, there are two associated complem-entation-symmetric Rote sequences $\mathbf{x}$ and $\mathbf{x}^{\prime}$ with $\mathbf{x}^{\prime}=E(\mathbf{x})$. The factors in $\mathbf{s}$ and its corresponding Rote sequences are closely related as shown below.
Proposition 49 ([85]; also see [68, Proposition 2] or [69, Lemma 2.7]).
Let $\mathbf{s}$ be a Sturmian sequence and let $\mathbf{x}$ be the complementation-symmetric Rote sequence such that $\mathbf{s}=\Delta(\mathbf{x})$. Then $u$ is a factor of $\mathbf{s}$ if and only if both words $v,v^{\prime}$ such that $u=\Delta(v)=\Delta(v^{\prime})$ are factors of $\mathbf{x}$. Furthermore, for every $n\geq 0$, $u$ occurs at position $n$ in $\mathbf{s}$ if and only if $v$ or $v^{\prime}$ occurs at position $n$ in $\mathbf{x}$.
@@
Let $\mathbf{x}$ be a complementation-symmetric Rote sequence. We clearly have $r_{\mathbf{x}}(0)=1$. Now, for $n\geq 1$[4, Theorem 8] states that $\operatorname{Pal}_{\mathbf{x}}(n)=2$. We finish the proof using Lemma 50 and Theorem 9(b). ∎
7 Rich reversal-closed sequences
Rich sequences have several equivalent definitions. It is known that a word $w$ contains at most $|w|+1$ palindromic factors [46]. A sequence is called rich if each factor contains the maximal number of palindromic factors.
@@ 7 Rich reversal-closed sequences
From [27, Theorem 1.1], the sequence $\mathbf{x}$ is rich if and only if the inequality in Theorem 17(b) is an equality. The result then follows from Theorem 9(b). ∎
Those among binary quasi-Sturmian sequences that are codings of rotations are rich, see [20, Theorem 19].
@@
Using Theorem 9(b) and [87, Corollaries 2.27 and 2.29], it is possible to bound the reflection complexity of rich sequences as follows.
@@
Other sequences have a reflection complexity satisfying the equality of Theorem 52. For instance, it is the case for complementation-symmetric sequences, sequences canonically associated with some specific Parry numbers, and sequences coding particular interval exchange transformations. For more details, see [14, Section 3].
@@ 8 Automatic sequences
In this section, we study the reflection complexity of automatic sequences. First, in a positional numeration system $U$ having an adder (i.e., addition is recognizable by an automaton reading $U$-representations), we show that if a sequence is $U$-automatic, then its reflection complexity is a $U$-regular sequence. Furthermore we show how to effectively compute a linear representation for the sequence, making use of the free software Walnut [74, 90]. Next, we explore the reflection complexity of some famous automatic sequences, namely the Thue–Morse, the period-doubling, generalized paperfolding, generalized Golay–Shapiro, and the Baum–Sweet sequences.
@@ 8.1 Reflection complexity is computably regular
We now show that the reflection complexity of an automatic sequence is regular.
@@
Let $U=(U(n))_{n\geq 0}$ be a positional numeration system such that there is an adder, and let $\mathbf{x}$ be a $U$-automatic sequence. Then $(r_{\mathbf{x}}(n))_{n\geq 0}$ is a $U$-regular sequence. Furthermore, a linear representation for $(r_{\mathbf{x}}(n))_{n\geq 0}$ is computable from the DFAO for $\mathbf{x}$.
@@
First we study the case of the Thue-Morse sequence. We want to emphasize the fact that the point of this example is not to reprove things that could be done more easily by appealing to existing theorems. It is to illustrate how the approach via Walnut can, in principle, carry out the various constructions for an arbitrary word, using Thue-Morse as an example simple enough where the matrices can actually be displayed.
We can compute a linear representation for the reflection complexity $r_{\mathbf{t}}(n)$ of the $2$-automatic Thue–Morse sequence $\mathbf{t}$, using the same approach as in the preceding section. Here we use the following Walnut code:
@@
    There is an automaton of 14 states that computes the first difference (-rt(+n1)rt(n))n0.
@@
    We can compute a linear representation for the first difference (-rt(+n1)rt(n))n0, and then use the “semigroup trick” [90, Section 4.11] to prove that the difference is bounded and find the automaton for it. It is displayed in Figure 2.
    Figure 2: Automaton computing (-rt(+n1)rt(n))n0 where t is the Thue–Morse sequence.
@@
A similar approach can be used to evaluate the reflection complexity for the period-doubling sequence p, which may be defined as the fixed point of the morphism 001 and 100. This gives us that p is 2-automatic as well. By analogy with Theorem 56, we can show that for all n0, we have =rp(+2n1)+ρp(n)1, and, for all n2, we have
@@
and we may similarly devise an analogue of part (c) of Theorem 56.
@@
Now, by a known result on the appearance function of paperfolding sequences [90, Theorem 12.2.1], we know that every length-14 factor of a paperfolding sequence pf appears in a prefix of length 109, which is in turn specified by the first 7 unfolding instructions. We can then simply examine each of the 56 length-14 factors of these 128 (finite) words and verify that no factor is reflected. ∎
A generalized Golay–Shapiro sequence g is defined by taking the running sum, modulo 2, of a paperfolding sequence pf. The famous Golay–Shapiro sequence (also called the Rudin–Shapiro sequence[56, 57, 86, 91] corresponds to the case of unfolding instructions 0(01)ω [4, Definition 6]. Note that a generalized Golay–Shapiro sequence is 2-automatic if and only if its corresponding generalized paperfolding sequence is 2-automatic.
The following analogue of Proposition 57 can be proved like Proposition 57, by appealing to a known result on the recurrence function of generalized Golay–Shapiro sequences [5, Proposition 4.1], giving that every length-15 factor of a paperfolding sequence pf appears in a prefix of length 2408, which is in turn specified by the first 12 unfolding instructions.
@@
    The reflection complexity of every generalized Golay–Shapiro sequence is the same, and takes the values 2, 3, 6, 10, 14, 22, 30, 42, 48, 62, 72, 83, 92, 103 for n[1,14].
@@
    The result for n15 follows from Item (a). For <n15 the result can be verified by enumeration of all length-2408 prefixes of paperfolding sequences specified by instructions of length 12.
@@
b=(b(n))n0=1101100101001001100100000100100101001001
be defined by =b(0)1 and for n1, b(n) is 1 if the base-2 expansion of n contains no block of successive zeros of odd length and 0 otherwise. It is 2-automatic as well. The factor complexity function for b starts with
@@
and the reflection complexity function for b with
@@
We can again compute a linear representation for (rb(n))n0 using the following Walnut code:
@@
Let b be the Baum–Sweet sequence. Then the first difference of the sequence rb(n) is 2-automatic, over the alphabet {1,2,,8}.
@@ 9 Further directions
The sequence c is not automatic in a given so-called addable numeration system (i.e., where there is an adder). Therefore, we cannot use Walnut, in this case. However, an inductive argument can be applied to prove that =rc(n)ρc(n) for all n13.
We propose the following natural questions.
Question 61.
@@
Question 63.
What is the reflection complexity of the Thue–Morse sequence over polynomial extractions, with regard to the work of Moshe [73]?
Question 64.
How can the upper bound in Theorem 21 be improved? If rx(n) is of the form Ω(n), then how can this be improved?
@@
How does the reflection complexity compare with other complexity functions, e.g., the complexity functions listed in Section 1?
This leads us to ask about the respective growths of the complexity functions listed in Section 1, in particular for morphic sequences. In this direction, recall that the factor complexity of a pure morphic sequence is either Θ(1), Θ(n), Θ(nloglogn), Θ(nlogn) or Θ(n2), see [76] (more details can be found, e.g., in [38]). The factor complexity of a morphic sequence is either Θ(n+1/1k) for some positive integer k, or else it is O(nlogn), see [45, 44]). As an illustration with a result that has not been already cited above, a comparison between growths for the factor complexity and the Lempel-Ziv complexity can be found in [39]. We end with an easy result for the growth of the reflection complexity in the case of pure morphic and morphic sequences.
@@
  • The reflection complexity of a pure morphic sequence is either Θ(1), Θ(n), Θ(nloglogn), Θ(nlogn) or Θ(n2).
  • The reflection complexity of a morphic sequence is either Θ(n+1/1k) for some positive integer k, or else it is O(nlogn).
@@
Use the inequalities in Theorem 9: for a sequence x and for all n0, we have 12ρx(n)rx(n)ρx(n). ∎
@@ Acknowledgments
We thank Boris Adamczewski for interesting discussions. John Campbell is grateful to acknowledge support from a Killam Postdoctoral Fellowship from the Killam Trusts, and thanks Karl Dilcher for a useful discussion. The research of Jeffrey Shallit is supported by NSERC grant 2024-03725. Manon Stipulanti is an FNRS Research Associate supported by the Research grant 1.C.104.24F. The authors are thankful for a number of detailed reviewer reports that have substantially improved this paper.